# through $\mathcal{R}$eal-$\mathcal{T}$ime $\mathcal{V}$ideo

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) increasingly excel at perception, understanding, and reasoning. However, current benchmarks inadequately evaluate their ability to perform these tasks continuously in dynamic, real-world environments. To bridge this gap, we introduce $\mathcal{RTV}\text{-}Bench$, a fine-grained benchmark for MLLM real-time video analysis. $\mathcal{RTV}\text{-}Bench$ includes three key principles: (1) Multi-Timestamp Question Answering (MTQA), where answers evolve with scene changes; (2) Hierarchical Question Structure, combining basic and advanced queries; and (3) Multi-dimensional Evaluation, assessing the ability of continuous perception, understanding, and reasoning. $\mathcal{RTV}\text{-}Bench$ contains 552 diverse videos (167.2 hours) and 4,631 high-quality QA pairs. We evaluated leading MLLMs, including proprietary (GPT-4o, Gemini 2.0), open-source offline (Qwen2.5-VL, VideoLLaMA3), and open-source real-time (VITA-1.5, InternLM-XComposer2.5-OmniLive) models. Experiment results show open-source real-time models largely outperform offline ones but still trail top proprietary models. Our analysis also reveals that larger model size or higher frame sampling rates do not significantly boost $\mathcal{RTV}\text{-}Bench$ performance, sometimes causing slight decreases. This underscores the need for better model architectures optimized for video stream processing and long sequences to advance real-time video analysis with MLLMs.
Figure 1: Specific examples of different tasks in RTV-Bench.
@@
represent the fundamental questions in the question group, and the answers are underlined.
is a dynamically answered question, and the correct answer is determined by the time of the query. The correct answer to the query will change at different times, and we have marked the correct answers that should be responded to at different query times.
@@ 1 Introduction
The ability to comprehend and respond to complex real-world scenarios in real-time represents a crucial challenge in the pursuit of general artificial intelligence [6, 24, 10]. Drawing inspiration from the remarkable achievements of large language models (LLMs) in a wide range of tasks [25, 11, 2], the potential of multimodal large language models (MLLMs) is increasingly being realized in the domain of visual scene analysis [48, 41, 5]. The research trajectory of Video-LLMs [21, 27, 4, 35, 33] is progressively advancing from the understanding of short, vision-centric videos [15, 13, 22, 19] to the integration of omni-modal information—including video, audio, and subtitles [3, 7, 17, 32, 18, 26]—for comprehensive long video analysis.
@@
Based on the above considerations, we introduce $\mathcal{RTV}\text{-}Bench$, featuring three core innovations designed to benchmark the continuous analysis capabilities—specifically perception, understanding, and reasoning—of MLLMs within real-time video contexts. First, the Multi-Timestamp Q&A mechanism challenges real-time tracking and state update by posing queries whose answers evolve within a video. Crucially, unlike benchmarks like OVO-Bench [20] that typically introduce different questions at different timestamps, RTV-Bench revisits the same conceptual query, where only the correct answer shifts as the scene unfolds. This approach more rigorously tests the model capacities for continuous analysis in real-time scenarios, surpassing single-query-single-answer and static evaluations. Second, the Hierarchical Question Structure enforces reliable, sequential reasoning by employing a basic-to-advanced design where higher-order questions logically depend on grasping foundational perceptions and understanding, thus mitigating cognitive shortcuts. Finally, Multidimensional Evaluation moves beyond single scores to provide granular diagnostics, assessing performance across eight fine-grained dimensions critical to continuous analysis in dynamic scenarios, offering a more insightful perspective on model capabilities and bottlenecks in handling real-time video information, detailed cases are shown in Figure 1.
Through RTV-Bench’s novel design and the detailed evaluations presented herein, we provide comprehensive insights into the current state of MLLMs for continuous video analysis: Results for all tested models on RTV-Bench reveal significant bottlenecks for existing MLLMs in real-time video analysis, as (1) most models achieve accuracy below 50%. Furthermore, (2) model performance shows minimal correlation with the number of input frames or model size. Notably, (3) models specifically designed for real-time video significantly outperform traditional offline video models, even the lowest-performing real-time model tested (VITA-1.5 [9]) surpasses a traditional offline counterpart (VideoLLaMA2 [4]). Building upon these findings, we further discuss promising research trajectories to enhance MLLM online video processing. We expect that our benchmark, analysis, and outlined future directions will contribute significantly to the development of more robust and capable online video MLLMs.
2 RTV-Bench
Figure 2: Video Categories and Distribution of Question Difficulty and Query Characteristics. (Left) RTV-Bench overs 3 key domains and 16 sub-class video types. (Center) Distribution of question difficulty levels across eight representative task types, measured by percentage-based performance ranges. (Right) Distribution of question queries by video length, categorized into Shallow, Moderate, and Deep levels. The bar heights indicate counts, while the line chart overlays query proportions for each duration bucket.
@@ 2.1 Challenge for Real-Time Video Analysis
Existing benchmarks for long video analysis [8, 38] attempt to gauge the true capacity of MLLMs by extending video duration, enhancing the difficulty of QA tasks, and introducing new sub-tasks. Other works [23] focus on real-time video scenarios, evaluating model responsiveness to queries in real-time contexts. However, current models face the risk of memory loss and attention shift, both in long and real-time video contexts. We categorize these challenges as failures in the continuous analysis capability of MLLMs. In video scenarios, MLLMs are considered to possess strong continuous analysis ability only if they can effectively recall prior visual information to accurately respond to user queries (similar to needle-in-a-haystack tasks in natural language settings) and maintain robust perception of incoming streams of new visual data and queries (akin to multimodal multi-turn dialogues).
@@ 2.2 Benchmark Overview
The RTV-Bench is designed to assess a model’s ability to perceive, understand, and reason in long video and real-time streaming contexts. Models should be adept at recognizing correct temporal query patterns. For example, during live sports queries (e.g., goalkeeper actions), models must prioritize current contexts over historical data. However, it should also retrieve relevant information from memory, such as identifying the goalkeeper from earlier footage. This necessitates that models not only possess exceptionally high fundamental scene comprehension capabilities, but also demonstrate effective understanding of all stages within the temporal flow—namely, comprehensive spatiotemporal comprehension encompassing the past, present, and future. To this end, we propose a series of sub-tasks for video continuous analysis, addressing both spatiotemporal elements within the video and the intrinsic capabilities of the model. The intrinsic model capabilities are divided into eight categories: Temporal Perception (TP), Scene Perception (SP), Visual Perception (VP), Future Prediction (FP), Phenomenological Understanding (PU), Intent Analysis (IA), Global Understanding (GU), and Spatiotemporal Reasoning (SR), see Figure 1 for details.
@@ 2.3 Benchmark Construction
The RTV-Bench dataset comprises $N=552$ videos sourced from the internet. A key characteristic is its question structure designed to directly probe temporal dynamics. For each video, questions are organized into sets. Each question set contains a minimum of three multiple-choice questions.
Crucially, the core design principle revolves around time-varying correct answers (Multi-Timestamp QA): within a set of questions, the same underlying conceptual query (for example, ’What is person A holding?’, ’Where is the car heading?’) is effectively evaluated at different points in time within the video. This means that the correct answer option changes depending on the specific timestamp or time interval referenced or implied by the question context or its associated earliest inferable timestamp (Figure 1). This challenges models not just to locate information, but to actively track changes and update their understanding of a situation as it evolves. This design directly targets the assessment of continuous understanding and the model’s sensitivity to temporal changes, see Figure 4 for cases.
@@
Table 1: Comparison of video QA benchmarks. MT: multi-timestamp questions. Labels: A (automatic), M (manual), A+M (both).
Benchmark MT Num. of QA Avg. Dur. (min) Total Dur. (h) Labels
Video-MME [8] 2.7k 17 254.0 M
MSRVTT [40] 73k 0.25 12.5 A
MSVD-QA [1] 13k 0.17 1.4 A
MovieChat-1k [31] 13k 9.4 156.0 M
MVBench [16] 4k 0.25 34.5 A+M
ActivityNet-QA [43] 58k 1.87 25.0 M
Vstream-Q [45] 3.5k 40 21.0
StreamBench [37] 1.8k 4.5 25.0 M
OVO-Bench [20] 2.8k 6 66.0 A+M
RTV-Bench (Ours) 4.6k 18 167.2 M
@@
Rigorous manual annotation by qualified experts was essential for RTV-Bench’s reliability in assessing continuous video analysis. While an LLM (Deepseek [25]) provided initial question templates, human annotators significantly refined questions to target dynamic scenes and temporal reasoning. They meticulously crafted questions where answers change over time and identified the earliest valid timestamp for each answer option in MTQA. This human-centric, multi-annotator process ensured robustness and created evaluations specifically focused on temporal dynamics in video.
@@
To ensure the benchmark’s quality, each video and its corresponding Q&A pairs underwent multiple rounds of review. We manually filtered videos based on length distribution and the presence of sub-scenes examining real-time event changes, resulting in high-quality videos focused on real-time analysis tasks. We conducted manual video-question alignment and precise timestamp checks, utilizing GPT-4 and human review to verify annotation format, translation, and sensitive information.
@@ 3.1 Experiment Setup
Our experiments were conducted across two A800 GPUs to comprehensively evaluate the performance of mainstream MLLMs on our benchmark. The selected models include VideoLLaMA2 [4], VideoLLaMA3 [44], GPT-4O [12], InternLM-XComposer2.5-OmniLive (IXC2.5-OL) [46], VITA1.5 [9], LLaVA-Video [21], LLaVA-OneVision [14], and Qwen2.5-VL [42]. These models were primarily tested with 7B to ensure a fair comparison under similar computational constraints.
@@
Table 2: Evaluation results on RTV-Bench. Perception, Understanding and Reasoning are different tasks. FQA notes that Foundational Video Question & Answersing without multi-timestamp answers. MTQA notes that Multi-Timestamp Question&Answersing with multi-timestamp answers.
@@ 3.2 Experiment Results
To investigate the effect of model scale, we evaluated Qwen2.5-VL across 3B-72B parameters. Figure 3 reveals non-monotonic scaling: medium-sized models (7B-34B) outperform both smaller and larger variants. All metrics fluctuate within a 5% range, with no configuration exceeding 34B’s maximums. This scale saturation implies current model architectures may lack the representational efficiency required for effective real-time processing, emphasizing the need for structural innovations rather than pure parameter growth.
@@
Figure 3 presents another counter-intuitive finding. While more input frames typically yield better performance in offline evaluations, our real-time setting shows no performance improvement with increased frames. In fact, IXC2.5-OL exhibits significant degradation. This suggests potential limitations in the model’s processing capacity or a breakdown mechanism, such as attention dilution. A potential limitation is that the current frame numbers may not adequately capture all critical moments or key information within long videos. Even with increased sampling density, crucial details might still be missed.
@@ 6 Conclusion
In this work, we introduce RTV-Bench, a Benchmark focuses on evaluating the video continuous analysis capabilities of MLLMs. We proposed relevant evaluation tasks and constructed RTV-Bench, which comprises 552 videos and 4,631 QA pairs. Extensive experiments conducted on diverse models reveal that: Models specifically designed for online video processing exhibit superior video continuous understanding capabilities, whereas simply increasing the number of uniformly sampled frames during training or inference stages currently shows limited effectiveness in enhancing performance on video continuous tasks.

### v4

# Shikang Wang6, Yixin Liu1, Hanbo Zhang7, Ying Ma1, Xuming Hu2,3

@@ Abstract
Multimodal Large Language Models (MLLMs) have made rapid progress in perception, understanding, and reasoning, yet existing benchmarks fall short in evaluating these abilities under continuous and dynamic real-world video streams. Such settings require models to maintain coherent understanding and reasoning as visual scenes evolve over time. We introduce $\mathcal{RTV}\text{-}Bench$, a fine-grained benchmark for real-time video analysis with MLLMs. It is built upon three key principles: multi-timestamp question answering, hierarchical question structures spanning perception and reasoning, and multi-dimensional evaluation of continuous perception, understanding, and reasoning. $\mathcal{RTV}\text{-}Bench$ comprises 552 diverse videos and 4,608 carefully curated QA pairs covering a wide range of dynamic scenarios. We evaluate a broad range of state-of-the-art MLLMs, including proprietary, open-source offline, and open-source real-time models. Our results show that real-time models generally outperform offline counterparts but still lag behind leading proprietary systems. While scaling model capacity generally yields performance gains, simply increasing the density of sampled input frames does not consistently translate into improved results. These observations suggest inherent limitations in current architectures when handling long-horizon video streams, underscoring the need for models explicitly designed for streaming video processing and analysis.
Figure 1: Representative examples illustrating the diverse task types evaluated in $\mathcal{RTV}\text{-}Bench$.
@@
denote fundamental questions within a question group, with their corresponding answers underlined.
indicates a dynamically answered question, where the correct response is determined by the query time. As the visual content evolves, the correct answer may change over time; we therefore annotate the appropriate answers corresponding to different query timestamps.
@@ 1 Introduction
The ability to comprehend and respond to complex real-world scenarios in real time remains a fundamental challenge in the pursuit of general artificial intelligence [6, 25, 10]. Motivated by the remarkable success of large language models (LLMs) across a broad spectrum of tasks [26, 11, 2], multimodal large language models (MLLMs) have recently emerged as a promising paradigm for visual scene understanding and reasoning [49, 42, 5]. In particular, the research trajectory of Video-LLMs [22, 28, 4, 36, 34] has evolved from early studies focused on short, vision-centric video clips [16, 14, 23, 20] toward more comprehensive modeling of long-form video content. An increasing body of work integrates omni-modal signals—including video, audio, and subtitles [3, 7, 18, 33, 19, 27]—to support richer contextual understanding and more robust long-horizon reasoning.
@@
Based on the above considerations, we introduce $\mathcal{RTV}\text{-}Bench$, featuring three core innovations designed to benchmark the continuous analysis capabilities—specifically perception, understanding, and reasoning—of MLLMs within real-time video contexts. First, the Multi-Timestamp Q&A mechanism challenges real-time tracking and state update by posing queries whose answers evolve within a video. Crucially, unlike benchmarks like OVO-Bench [21] that typically introduce different questions at different timestamps, RTV-Bench revisits the same conceptual query, where only the correct answer shifts as the scene unfolds. This approach more rigorously tests the model capacities for continuous analysis in real-time scenarios, surpassing single-query-single-answer and static evaluations. Second, the Hierarchical Question Structure enforces reliable, sequential reasoning by employing a basic-to-advanced design where higher-order questions logically depend on grasping foundational perceptions and understanding, thus mitigating cognitive shortcuts. Finally, Multidimensional Evaluation moves beyond aggregate scores to provide fine-grained diagnostic insights, assessing model performance across eight dimensions that are critical for continuous analysis in dynamic scenarios. This evaluation offers a more informative view of model capabilities and limitations in real-time video understanding; detailed examples are presented in Figure 1.
Through the novel design of $\mathcal{RTV}\text{-}Bench$ and the comprehensive evaluations presented in this work, we provide systematic insights into the current state of MLLMs for continuous video analysis. Results across all evaluated models reveal substantial bottlenecks in real-time video understanding: ❶ most models achieve accuracies below 50%; ❷ overall performance shows a clear positive correlation with model scale, whereas increasing the number of input frames yields only marginal and non-monotonic gains; and ❸ models explicitly designed for streaming video processing consistently outperform traditional offline video models. Notably, even the lowest-performing real-time model evaluated, VITA-1.5 [9], surpasses a representative offline counterpart, VideoLLaMA2 [4]. Building on these findings, we further discuss promising research directions for online video analysis with MLLMs.
2 Real-Time Video Understanding for MLLMs: RTV-Bench
Figure 2: Video categories and distributions of question difficulty and query characteristics. (Left) RTV-Bench covers three key domains and 16 video subcategories. (Center) Distribution of question difficulty levels across eight representative task types, quantified by percentage-based performance ranges. (Right) Distribution of question queries with respect to video length, categorized into Shallow, Moderate, and Deep levels. Bar heights indicate query counts, while the overlaid line chart shows the proportion of queries within each duration bucket.
@@ 2.1 Challenge for Real-Time Video Analysis
Existing benchmarks for long video analysis [8, 39] attempt to gauge the true capacity of MLLMs by extending video duration, enhancing the difficulty of QA tasks, and introducing new sub-tasks. Other works [24] focus on real-time video scenarios, evaluating model responsiveness to queries in real-time contexts. However, current models face the risk of memory loss and attention shift, both in long and real-time video contexts. We categorize these challenges as failures in the continuous analysis capability of MLLMs. In video scenarios, MLLMs are considered to possess strong continuous analysis ability only if they can effectively recall prior visual information to accurately respond to user queries (similar to needle-in-a-haystack tasks in natural language settings) and maintain robust perception of streams of new visual data and queries (akin to multimodal multi-turn dialogues).
@@ 2.2 Benchmark Overview
The RTV-Bench is designed to assess a model’s ability to perceive, understand, and reason in long video and real-time streaming contexts. Models should be adept at recognizing correct temporal query patterns. For example, during live sports queries (e.g., goalkeeper actions), models must prioritize current contexts over historical data. However, it should also retrieve relevant information from memory, such as identifying the goalkeeper from earlier footage. This necessitates that models not only possess exceptionally high fundamental scene comprehension capabilities, but also demonstrate effective understanding of all stages within the temporal flow—namely, comprehensive spatiotemporal comprehension encompassing the past, present, and future. To this end, we propose a series of sub-tasks for video continuous analysis, addressing both spatiotemporal elements within the video and the intrinsic capabilities of the model. The intrinsic model capabilities are divided into eight categories: Temporal Perception (TP), Scene Perception (SP), Visual Perception (VP), Future Prediction (FP), Phenomenological Understanding (PU), Intent Analysis (IA), Global Understanding (GU), and Spatiotemporal Reasoning (SR). An overview of these task categories is provided in Figure 1.
@@ 2.3 Benchmark Construction
The RTV-Bench dataset comprises 552 videos sourced from the internet. A key characteristic is its question structure designed to directly probe temporal dynamics. For each video, questions are organized into sets. Each question set contains about three multiple-choice questions.
Crucially, the core design principle centers on time-varying correct answers through Multi-Timestamp QA. Specifically, within a set of questions, the same underlying conceptual query (e.g., “What is person A holding?” or “Where is the car heading?”) is evaluated at multiple points in time throughout the video. As a result, the correct answer may differ depending on the specific timestamp or temporal interval referenced—or implicitly required—by the question context or its earliest inferable timestamp (Figure 1). Rather than merely locating relevant information, models are therefore required to actively track temporal changes and continuously update their understanding as the scene evolves. This design directly targets the evaluation of continuous temporal understanding and a model’s sensitivity to dynamic state transitions; representative examples are shown in Figure 4.
@@
Table 1: Comparison of video QA benchmarks. MT denotes multi-timestamp questions. Labels: A (automatic), M (manual), A+M (both).
Benchmark MT #QA (k) Avg. Duration (min) Total Duration (h) Labels
Video-MME [8] $2.7$ $17.00$ $254.0$ M
MSRVTT [41] $73.0$ $0.25$ $12.5$ A
MSVD-QA [1] $13.0$ $0.17$ $1.4$ A
MovieChat-1k [32] $13.0$ $9.40$ $156.0$ M
MVBench [17] $4.0$ $0.25$ $34.5$ A+M
ActivityNet-QA [44] $58.0$ $1.87$ $25.0$ M
Vstream-Q [46] $3.5$ $40.00$ $21.0$
StreamBench [38] $1.8$ $4.50$ $25.0$ M
OVO-Bench [21] $2.8$ $6.00$ $66.0$ A+M
OVBench [12] $7.0$ A
RTV-Bench (Ours) $4.6$ $18.00$ $167.2$ M
@@
Rigorous manual annotation by qualified experts underpins the reliability of RTV-Bench for evaluating continuous video understanding. We leverage an LLM (DeepSeek [26]) only to produce initial question templates, and human annotators then refine every question to better reflect dynamic scenes and the demands of temporal reasoning. In particular, annotators intentionally craft questions whose correct answers evolve over time and systematically determine the earliest valid timestamp associated with each answer option in the MTQA setting. This human-centric, multi-annotator protocol strengthens annotation robustness and enables RTV-Bench to explicitly evaluate models’ sensitivity to temporal dynamics in video.
@@
To ensure the benchmark’s quality, each video and its corresponding Q&A pairs underwent multiple rounds of review. We manually filtered videos based on length distribution and the presence of sub-scenes examining real-time event changes, resulting in high-quality videos focused on real-time analysis tasks. We conducted manual video-question alignment and precise timestamp checks, utilizing GPT-4 and human review to verify annotation format and sensitive information.
@@ 3.1 Experiment Setup
Our experiments were conducted on two NVIDIA A800 GPUs to comprehensively evaluate the performance of mainstream multimodal large language models (MLLMs) on our benchmark. We consider a diverse set of representative models, including VideoLLaMA2 [4], VideoLLaMA3 [45], GPT-4o [13], InternLM-XComposer2.5-OmniLive (IXC2.5-OL) [47], VITA1.5 [9], VideoChat-Online (4B) [12], LLaVA-Video [22], LLaVA-OneVision [15], and Qwen2.5-VL [43].
To ensure a fair comparison under comparable computational budgets, most models are evaluated using configurations around the 7B scale when available, while smaller models (e.g., VideoChat-Online at 4B) are evaluated at their native parameter size. All models are tested under a unified uniform frame sampling protocol. Specifically, for models that support variable frame inputs (e.g., Qwen2.5-VL), we evaluate multiple sampling settings with 8, 16, 32, and 64 uniformly sampled frames. For each model, we report in the main tables the best-performing configuration across different frame counts, following the same evaluation protocol for all baselines. Other models are evaluated analogously using their supported frame sampling ranges, and their strongest results are reported for comparison.
@@
Table 2: Evaluation results on RTV-Bench. Perception, Understanding, and Reasoning denote different task categories. FQA refers to foundational video question answering without multi-timestamp supervision. MTQA refers to multi-timestamp question answering with time-varying correct answers. Scores are computed using group-aware Q2 evaluation.
@@ 3.2 Experiment Results
We analyze the effect of model scale by evaluating Qwen2.5-VL from 3B to 72B parameters under different frame sampling budgets (8–64 frames), as shown in Figure 3(c). Overall accuracy exhibits a clear and largely monotonic improvement with increasing model size. Specifically, the 72B model consistently achieves the highest performance across all frame settings, reaching up to 40.78% with 64 frames, while smaller models (3B–32B) remain below 40%.
Notably, scaling benefits are consistent but moderate, with absolute gains from 3B to 72B on the order of $\sim$2–3 points, suggesting diminishing returns at larger scales. In addition, model scale interacts with temporal resolution: larger models benefit more reliably from increased frame counts, whereas smaller models show non-uniform or even fluctuating trends when additional frames are introduced. These observations indicate that while parameter scaling remains beneficial for real-time video understanding, its effectiveness is increasingly constrained by architectural and temporal modeling capacities, highlighting the importance of improving temporal representation efficiency beyond naive model enlargement.
@@
Figure 3(c,d) jointly examine the effect of frame sampling density across model scales and architectures. From Figure 3(c), increasing the number of frames from 8 to 64 does not yield consistent accuracy gains across model sizes. While larger models (e.g., 72B) show modest improvements with more frames, smaller and medium-scale models exhibit non-monotonic or saturated trends, indicating limited benefit from denser temporal sampling. In some cases (e.g., 3B and 32B), additional frames lead to marginal gains or even slight regressions.
This phenomenon becomes more pronounced in Figure 3(d), which aggregates performance across models. Average accuracy remains largely stable as frame count increases, while the total score—reflecting global understanding—often declines. Notably, IXC2.5-OL suffers a clear performance drop with more frames, suggesting that excessive temporal inputs may overwhelm the model’s effective processing capacity. Together, these results indicate that simply increasing frame numbers is insufficient for improving real-time video understanding and may instead introduce redundancy or attention dilution. This highlights the need for temporally selective, adaptive frame utilization strategies rather than uniform increases in sampling density.
@@ 6 Conclusion
In this work, we introduce RTV-Bench, a benchmark designed to systematically evaluate the continuous video understanding and real-time reasoning capabilities of multimodal large language models (MLLMs). RTV-Bench comprises 552 long-form and streaming-style videos paired with 4,608 carefully constructed QA instances, targeting time-varying perception, understanding, and reasoning under realistic online settings.
Extensive experiments across a diverse set of models reveal two key findings. First, models explicitly designed for online or streaming video processing consistently demonstrate stronger continuous understanding capabilities than general-purpose counterparts. Second, both parameter scaling and uniformly increasing the number of sampled frames during training or inference yield only limited and sometimes inconsistent performance gains. In particular, denser temporal sampling often leads to performance saturation or degradation, highlighting inherent limitations in current architectures when handling long or high-density visual streams. These observations underscore that effective real-time video understanding requires principled temporal modeling and selective information aggregation, rather than relying on naive increases in model size or frame counts.